OpenStudy (anonymous):

HELP IN 6th grade Solve for t. t – 15.7 = 53.4 A. 69.1 B. 68.1 C. 38.7 D. 37.7

OpenStudy (shamil98):

add 15.7 to both sides to get the value of t.

OpenStudy (anonymous):

d

OpenStudy (shamil98):

@melancholymushroom , that's false..

OpenStudy (shamil98):

anyways, don't just give out the answer, if possible give an explanation or guide them etc

OpenStudy (anonymous):

I THINK IT IS B

OpenStudy (anonymous):

a my bad

OpenStudy (charlotte123):

t – 15.7 = 53.4 + 15.7 + 15.7 53.4 + 15.7 = t <----

OpenStudy (shamil98):

Close. what is 53.4 + 15.7 ?

OpenStudy (anonymous):

its a

OpenStudy (charlotte123):

@shamil98 He he he \(\huge\color{black}{\frak{Wrong~Timing~Mortal}}\)

OpenStudy (anonymous):

IT IS A 69.1 I KNOW am i right

OpenStudy (charlotte123):

@angelv00 Yep :)

OpenStudy (anonymous):

YAY THANKS

OpenStudy (charlotte123):

No problem! Take care!

OpenStudy (anonymous):

YOU TOO

Directrix (directrix):

@melancholymushroom Code of conduct: OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok i get it